# Soft-TransFormers for Continual Learning

### v2

@@ 5 Experiments
We validate our Soft-TF on several benchmark datasets against continuous learning baselines in (1) Class-Incremental Learning (CIL) based on ViT models  [^1] https://github.com/ihaeyong/Soft-TF.git and (2) CIL based on language models  [^2] https://github.com/ihaeyong/LLM-Soft-TF.git.
@@ 5.1 Experimental Settings
Datasets. (1) CL based on ViT models: we evaluate our method mainly on 1) 10/20-Split-CIFAR100 (Krizhevsky et al., 2009), constructed by splitting the 100 classes into 10 tasks/20 tasks. 2) 10-Split-TinyImageNet (Abai and Rajmalwar, 2019), constructed by splitting the 200 classes into 10 tasks. 3) 10-Split-ImageNet-R (Hendrycks et al., 2021), constructed by splitting the 200 classes into 10 tasks. To show our effectiveness, we additionally compare our method with the baselines on 5-Split-CUB200 and 10-Split-TinyImageNet. (2) CL based on language models: we test our method on the widely adopted continual learning benchmarks for language models following de Masson D’Autume et al. (2019), which use five text classification datasets (Chen et al., 2020; Zhang et al., 2015) including AG News (news classification), Yelp (sentiment analysis), DBPedia (Wikipedia article classification), Amazon (sentiment analysis) and Yahoo Answers (Q&A classification).
Baselines. To validate the powerfulness of our method in (1) CL based on ViT models, we compare our results with various CIL baselines including, L2P (Wang et al., 2022c), DualPrompt (Wang et al., 2022b), and DualPrompt-PGP (Qiao et al., 2024). Moreover, to show the effectiveness of our method in (2) CL based on language models, we set MBPA++ (de Masson D’Autume et al., 2019), IDBR (Huang et al., 2021), SLM (Bohao et al., 2024), and DualPrompt (Wang et al., 2022b) as baselines.
Table 2: Performances of Language Class Incremental Learning (CIL, Bert) in terms of Accuracy, Forgetting, Trainable parameters, and Time complexity on Bert benchmark datasets. “DR”: data replay. $\ast$ denotes our reproduced results.
Order: 4 Order: 5 Order: 6 Order: 7 Average Bert-base (110.0M)
Method DR Task ID ACC/Forget ACC/Forget ACC/Forget ACC/Forget ACC/Forget #\#Tr.Params Tr./Test[sec]
MBPA++ $\checkmark$ L.Adapter 74.9 / 3.71 73.1 / 2.76 74.9 / 2.89 74.1 / 3.00 74.3 / 3.09 110.0M 30.12K / 87.43
IDBR $\checkmark$ L.Adapter 75.9 / 3.46 76.2 / 2.33 76.4 / 2.88 76.7 / 2.90 76.3 / 2.89 110.0M 27.73K / 86.22
SLM Retrieval 78.2 / 0.70 77.8 / 0.70 78.0 / 0.70 78.2 / 0.70 78.1 / 0.70 7.32M 17.51K / 73.36
DualPrompt-L[3,4,5] Prompt ID 69.7 / 0.02 67.5 / 0.04 69.1 / 0.04 68.8 / 0.03 68.9 / 0.03 0.24M 13.86K / 70.14
DualPrompt-L[10,11,12] Prompt ID 67.2 / 0.06 65.1 / 0.09 66.5 / 0.08 66.3 / 0.07 66.3 / 0.08 0.24M 13.86K / 70.14
+ Adapter-L[all], r=16 Prompt ID 75.0 / 0.00 75.1 / 0.00 75.1 / 0.00 75.1 / 0.00 75.1 / 0.00 7.54M 19.45K / 75.34
+ LoRA-L[all], r=96 Prompt ID 78.9 / 0.00 78.9 / 0.00 78.9 / 0.00 78.9 / 0.00 78.9 / 0.00 7.28M 21.57K / 76.61
+ Soft-TF-L[3,4,5] (SOTA) Prompt ID 79.3 / 0.00 79.6 / 0.00 79.4 / 0.00 79.4 / 0.00 79.4 / 0.00 7.23M 15.51K / 71.12
+ Soft-TF-L[10,11,12] Prompt ID 79.5 / 0.00 79.3 / 0.00 79.3 / 0.00 79.5 / 0.00 79.4 / 0.00 7.23M 15.53K / 71.12
[…]
Table 3: Performances of Vision Class Incremental Learning (CIL, ViTs) in terms of Pretained-dataset (ImageNet-21K, SAM, DINO) and Task-IDs on 10-Split-CIFAR100 and 5-Split-CUB200.
10-Split-CIFAR100 5-Split-CUB200
Method Pretrained-dataset Task ID ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$)
DualPrompt ImageNet-21K Prompt ID 86.50 5.77 82.02 4.23
DualPrompt-PGP ImageNet-21K Prompt ID 86.92 5.35 82.46 3.76
DualPrompt SAM Prompt ID 86.11 6.08 82.02 4.73
DualPrompt DINO Prompt ID 64.18 23.81 50.88 10.10
 +Soft-TF ImageNet-21K Prompt ID 92.42 2.44 76.17 9.04
 +Soft-TF (SOTA) ImageNet-21K Gradient ID 97.87 0.21 87.93 0.66
 +Soft-TF (SOTA) SAM Gradient ID 97.87 0.21 87.93 0.66
 +Soft-TF DINO Gradient ID 84.50 12.27 69.79 10.93
Upper-Bound of Soft-TF - - 93.90 - 85.56 -
@@
(a) L2P-PGP v.s. Soft-Transformer (b) DualPrompt v.s. Soft-Transformer
Figure 5: Layer-wise(L[$\ast$]) Performances of Soft-TF on 10-Split-CIFAR100. Note that L[9,10,11] denotes Soft-TransFormer of 9, 10, 11 Layers.
@@ 5.2 Performances
Performances of Soft-TF on Vision CIL. We compare our Soft-TransFormers (Soft-TF) with state-of-the-art CIL baselines, as shown in Table 1. Our Soft-TF significantly outperforms all baselines: Adapter and LoRA and upper-bounds of Soft-TF, including L2P and DualPrompt, in accuracy and forgetting measurements. The time complexity of training and testing does not seem to be a significant issue under the same trainable parameters. The performance gain of Soft-TF is especially notable in DualPrompt-based learning compared to L2P, suggesting the importance of global prompt-tuning and multi-head attention prompt-tuning in DualPrompt. Additionally, task-identity inference using Gradient-ID is crucial for achieving full fine-tuning results in CIL.
Performances of Soft-TF on Language CIL. Our Soft-TF significantly outperforms all baselines including MBPA++, IDBR, SLM, and DualPrompt in accuracy and forgetting measurements, as shown in Table 2. Replay-based methods, i.e., MBPA++ and IDBR, carry a computational burden since they replay the previous task’s samples and fine-tune 110.0M parameters. SLM shows inefficient time complexity since it takes two stages (warmup and fine-tuning) in the training step and the retrieval process in the inference step. Moreover, Soft-TF performs better than LLM-based fine-tuning baselines such as Adapter and LoRA under the same trainable parameter settings. Soft-TF trains and tests faster than LLM-based baselines since Soft-TF fine-tunes sparse layers (i.e., [3,4,5] layers). Note that LoRA failed to fine-tune onto sparse [3,4,5] layers.
5.3 Ablation Studies
Well-initialized LTH (WLTH) on Vision CIL. To demonstrate the efficacy of our proposed method on Well-initialized Lottery Ticket Hypothesis (WLTH) backbones, we evaluate our Soft-TransFormers (Soft-TF) by extending two distinct pre-trained models, ViT-DINO and ViT-SAM (Caron et al., 2021; Chen et al., 2021). As shown in Table 3, we tested our method on the 10-Split-CIFAR100 and 5-Split-CUB200 datasets using three pre-trained ViTs: ImageNet-21K, DINO, and SAM, further validating the effectiveness of our approach on non-ImageNet datasets (Krizhevsky et al., 2009; Wah et al., 2011). Surprisingly, when initialized with ImageNet-21K and SAM, DualPrompt-Soft-TF with ImageNet-21K achieved the same performance levels. Moreover, DualPrompt-Soft-TF outperformed all baselines, i.e., DualPrompt-PGP, on both benchmark datasets, indicating that well-initialized weights provide better generalization in continual learning scenarios.
@@ 5.3 Ablation Studies
Random initialization. Random initialization of Soft-Transformer’s weights plays a critical role when leveraging well-pretrained models like Vision Transformers (ViTs). The optimal training point is the parameters of a well-pretrained model. Among the initialization methods, Uniform initialization for Soft-TransFormer satisfies this requirement effectively. To validate these claims, we analyze the impact of common random initialization methods, including Xavier, Kaiming, Normal, and Uniform Initialization, as shown in Table 4. The results demonstrate that the same well-initialization point leads to independent optimal task performance, particularly with Gradient ID inference. Furthermore, this ablation study strengthens our Soft-TF with state-of-the-art-perfomances inspired by the Well-initialized Lottery Ticket Hypothesis (WLTH).
@@
Table 4: Random initialized Performances of Class Incremental Learning (CIL, ViTs) in terms of accuracy and forgetting on 10-Split-CIFAR100. Note ”w/o FF” denotes ”Soft fine-tuning without FeedForward (FF)” networks. DualPrompt’s parameters are pretrained on ImageNet-21K.
@@
DualPrompt Task ID Random Init. ACC($\uparrow$) Forget($\downarrow$)
 +Soft-TF-L[10,11,12] w/o FF Prompt ID Xavier 90.59 3.85
 +Soft-TF-L[10,11,12] w/o FF Prompt ID Kaiming 90.72 3.63
 +Soft-TF-L[10,11,12] w/o FF Prompt ID Normal 90.45 3.78
 +Soft-TF-L[10,11,12] w/o FF Prompt ID Uniform(1.0, 1.0) 92.35 2.98
 +Soft-TF-L[10,11,12] w/o FF Gradient ID Uniform(1.0, 1.0) 98.05 0.25
Upper-Bound of Soft-TF 93.90 -
Layer-wise Inspections. We analyze the layer-wise performance of Soft-Transformer with respect to L2P and DualPrompt on the 10-Split-CIFAR100 dataset to identify the optimal configurations, as shown in Figure 5. Our observations reveal that the global prompt in DualPrompt influences Soft-Transformer’s performance differently in L2P and DualPrompt settings. In L2P-PGP, the best performance was achieved with Soft-TransFormers applied to the lower layers ((a) L2P-Soft-TF-L[1,2]-PGP), whereas in DualPrompt, the higher layers ((b) DualPrompt-Soft-TF-L[10,11,12]) yielded the best results. Notably, DualPrompt-Soft-TF-L[10,11,12] without PGP demonstrated impressive performance, achieving almost zero forgetting (0.21). These findings suggest that our approach could significantly enhance the effectiveness of large-scale Transformer models in continual learning scenarios.
@@
Attention Layer. We compare attention maps from the 12th attention layer (L[12]) of two different models, DualPrompt and Soft-TF, applied to category 7 of the 10-Split-CIFAR100 dataset, as shown in Figure 6. While DualPrompt distributes its attention more uniformly, Soft-TF prioritizes specific patches of the input image, likely those more relevant to Category 7. These attention maps highlight the differences in how DualPrompt and Soft-TF allocate attention in a Vision Transformer. Soft-TF’s sharper focus on essential features aligns with the reported superior performance on the 10-Split-CIFAR100 dataset. This visualization emphasizes the interpretability and task-specific optimization of Soft-TF over the more generalized behavior of DualPrompt.
@@ 6 Conclusion
Inspired by Well-initialized Lottery Ticket Hypothesis (WLTH), which provides suboptimal fine-tuning solutions, we proposed a novel fully fine-tuned continual learning (CL) method referred to as Soft-TransFormers (Soft-TF). Soft-TF sequentially learns and selects an optimal soft-network for each task. During sequential training in CL, an well-initialized Soft-TF mask optimizes the weights of sparse layers to obtain task-adaptive soft (real-valued) networks, while keeping the well-pre-trained layer parameters frozen. In inference, the identified task-adaptive network of Soft-TF masks the parameters of the pre-trained network, mapping to an optimal solution for each task and minimizing Catastrophic Forgetting (CF) - the soft-masking preserves the knowledge of the pre-trained network. Extensive experiments on the Vision Transformer (ViT) and the Language Transformer (Bert) demonstrated the effectiveness of Soft-TF, achieving state-of-the-art performance across Vision and Language CIL scenarios.
@@ A.1 Analysis of Soft-TransFormers (Soft-TF)
where $\bm{w}_{m}\neq\bm{w}_{q}$ since $\bm{m}$ is learnable parameters in Soft-TransFormers. Specifically, we could assume that $\bm{w}_{m}=(\bm{w}^{Q}\odot\bm{m}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{m}^{K})^{T}$ and $\bm{w}_{p}=(\bm{w}^{Q}\odot\bm{1}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{1}^{K})^{T}$ of Equation 6.
Theorem 2.
@@
Proof  Using algebraic manipulations (completing the square), we obtain:
@@
In terms of Soft-TransFormers $\bm{w}_{m}=(\bm{w}^{Q}\odot\bm{m}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{m}^{K})^{T}$ of Equation 6, we have
@@
Refer to caption Refer to caption
(a) DualPrompt-PGP v.s. Soft-TF (b) DualPrompt-WSN v.s. Soft-TF

### v3

@@ 5 Experiments
We validate our Soft-TF on several benchmark datasets against continuous learning baselines in Class-Incremental Learning (CIL).
@@ 5.1 Experimental Settings
Datasets.
We evaluate our method mainly on 1) 10/20-Split-CIFAR100 (Krizhevsky et al., 2009), constructed by splitting the 100 classes into 10 tasks/20 tasks. 2) 10-Split-TinyImageNet (Abai and Rajmalwar, 2019), constructed by splitting the 200 classes into 10 tasks. 3) 10-Split-ImageNet-R (Hendrycks et al., 2021), constructed by splitting the 200 classes into 10 tasks. To show our effectiveness, we additionally compare our method with the baselines on 5-Split-CUB200.
Baselines. To validate the powerfulness of our method, we compare our results with various CIL baselines including, L2P (Wang et al., 2022d), DualPrompt (Wang et al., 2022c), DualPrompt-PGP (Qiao et al., 2024), and NoRGa (Le et al., 2024). At inference time, the system determines the task identity for an arbitrary test sample $\bm{x}$ to select the appropriate task-specific nuances and ultimately achieve full fine-tuning performance. This is accomplished by using the Prompt ID mechanism, which simply chooses the best-matched task index $t$ by finding the minimum value of a function $\gamma$ comparing the query representation of the sample, $q(\bm{x})$, against all learned task keys $\bm{k}_{t}$, expressed as $\text{argmin}_{t}\gamma(q(\bm{x}),\bm{k}_{t})$.
Table 2: Performances of DualPrompt+Soft-TF in terms of ViTs pretained-dataset (ImageNet-21K) on 10-Split-ImageNet-R.
DualPrompt 10-Split-ImageNet-R
+Soft-TF #Tr.Prams ACC($\uparrow$) Forget($\downarrow$) Tr./Test[sec]
ViT-Ti/16 (5.7M) 0.41M 49.34 9.02 15.10K / 50
ViT-S/16 (22M) 1.71M 65.12 6.67 15.22K / 51
ViT-B/16 (86M) 6.93M 76.62 5.30 15.35K / 52
ViT-L/16 (307M) 12.50M 78.86 4.46 28.57K / 99
@@
(a) Comparisons with baselines (b) Layer-wise Performances
Figure 4: Layer-wise(L[$\ast$]) Performances of DualPrompt+Soft-TF on 10-Split-CIFAR100. Note that L[10,11,12] denotes Soft-Transformer of 10, 11, 12 Layers.
@@ 5.2 Performances
Performances of Soft-TF. We compare our Soft-TF with state-of-the-art CIL baselines, as shown in Table˜1. Soft-TF consistently outperforms prompt-based and parameter-efficient fine-tuning methods—including L2P, DualPrompt, NoRGa, Adapter, LoRA, and their corresponding upper-bounds—across accuracy, forgetting, and efficiency metrics. In particular, Soft-TF exhibits substantial improvements over NoRGa, a strong prompt-based baseline that mitigates task interference, yet remains limited by its reliance on global or head-level prompt modulation. By contrast, Soft-TF directly modulates attention projection matrices through task-specific soft subnetworks, enabling fine-grained and continuous adaptation within a frozen Transformer backbone, which results in higher accuracy and significantly reduced forgetting across all benchmarks. Despite using comparable trainable parameter budgets, Soft-TF achieves these gains without additional training or inference cost. Remarkably, Soft-TF even surpasses the Upper-Bound (Joint) of DualPrompt using prompt baselines: DualPrompt and NoRGa on all evaluated benchmarks, as it explicitly identifies and optimizes distinct task-specific soft subnetworks for each incremental task, achieving near-optimal per-task specialization, whereas joint training is constrained to a single shared solution that must compromise across tasks.
Comparisions with baselines. Figure˜4 of (a) shows that Soft-TF consistently outperforms established parameter-efficient tuning baselines - including PGP, Adapter, and LoRA - across all sessions in 10-Split CIFAR-100, demonstrating superior resistance to task interference and better retention of previously learned knowledge. As shown in Figure˜4 of (b) the layer-wise analysis, even when fine-tuning only selected layers (e.g., layers 10, 11, 12), Soft-TF exhibits a range of performance behaviors depending on the chosen layer configuration, yet it still maintains a clear advantage over LoRA in every setting. This indicates that Soft-TF not only provides stronger overall continual learning stability but also preserves its effectiveness under various layer-wise adaptation strategies, highlighting its robustness and versatility.
5.3 Variants of ViTs.
Across different ViT backbones, Soft-TF exhibits consistent advantages in both parameter efficiency and continual learning stability on 10-Split-ImageNet-R, as shown in Table˜2. The lightweight ViT-Ti/16 uses only 0.41M trainable parameters but yields lower accuracy (49.34%) and higher forgetting (9.02), whereas scaling up to ViT-S/16, ViT-B/16, and ViT-L/16 increases trainable parameters to 1.71M, 6.93M, and 12.50M, respectively, resulting in markedly improved performance. In particular, Soft-TF with ViT-L/16 achieves the best overall results - 78.86% accuracy and only 4.46 forgetting - while maintaining comparable training and test times ($\approx$28K/99 sec). Figure˜5 shows that larger backbones maintain higher average accuracy across all 10 sessions, with ViT-L/16 degrading the slowest and sustaining a clear performance margin over ViT-B/16, ViT-S/16, and ViT-Ti/16. Overall performance differences are further influenced by each backbone’s representational capacity and initialization quality, which shape how effectively Soft-TF can adapt while resisting catastrophic forgetting. Moreover, while increasing the number of trainable parameters generally leads to better continual learning performance, this improvement naturally comes with higher computational cost, reflecting the typical trade-off between model capacity and efficiency.
Figure 5: Performances of DualPrompt+Soft-TF: in terms of ViT models on 10-Split-ImageNet-R. The performance of Soft-TF-ViT-B/16 is comparable to that of Soft-TF-ViT-L/16.
5.4 Well-initialized Soft-TF
@@
Random initialization. Random initialization of Soft-Transformer’s weights plays a critical role when leveraging well-pretrained models like Vision Transformers (ViTs). The optimal training point is the parameters of a well-pretrained model. Among the initialization methods, Uniform initialization for Soft-TransFormer satisfies this requirement effectively. To validate these claims, we analyze the impact of common random initialization methods, including Xavier, Kaiming, Normal, and Uniform Initialization, as shown in Table˜3. The results demonstrate that the same well-initialization point leads to independent optimal task performance. Furthermore, this ablation study strengthens our Soft-TF with state-of-the-art-perfomances inspired by the Well-initialized Lottery Ticket Hypothesis (WLTH).
Well-initialized LTH (WLTH). Based on Table˜4, we evaluated our Soft-TF on the 10-Split-CIFAR100 and 5-Split-CUB200 datasets using three distinct pre-trained ViT models (ImageNet-21K, SAM, and DINO) to demonstrate its efficacy, including on Well-initialized Lottery Ticket Hypothesis (WLTH) backbones. For the 10-Split-CIFAR100 benchmark, the ImageNet-21K backbone achieved the highest accuracy (92.35%) and the lowest forgetting rate (2.98%), closely followed by the SAM backbone (92.30% ACC, 2.99% Forget). Similarly, for the 5-Split-CUB200 benchmark, the ImageNet-21K backbone led with 76.17% ACC and 9.04% Forget, which was also nearly matched by the SAM backbone (76.11% ACC, 9.01% Forget). In contrast, the DINO backbone consistently showed the lowest performance across both datasets. These trends are consistent with our theoretical analysis, which predicts faster convergence and bounded parameter drift when optimization is initialized near a well-pretrained solution and confined to a low-dimensional soft-parameter space. These results confirm the effectiveness of Soft-TF and suggest that using well-initialized weights provides superior generalization and performance in continual learning settings.
@@
Table 3: Random initialized Performances of DualPrompt+Soft-TF on 10-Split-CIFAR100. Note DualPrompt’s parameters are pretrained on ImageNet-21K.
@@
Random Init. ACC($\uparrow$) Forget($\downarrow$)
Xavier 90.59 3.85
Kaiming 90.72 3.63
Normal 90.45 3.78
Uniform(1.0, 1.0) 92.35 2.98
Table 4: Performances of Soft-TF in terms of Pretained-dataset (ImageNet-21K, SAM, DINO) on 10-Split-CIFAR100 and 5-Split-CUB200.
DualPrompt 10-Split-CIFAR100 5-Split-CUB200
+Soft-TF ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$)
ImageNet-21K 92.35 2.98 76.17 9.04
SAM 92.30 2.99 76.11 9.01
DINO 84.50 12.27 69.79 10.93
5.5 Soft-MSA Layers
The Soft-TF originates from learned parameters distributed with $\mu\approx 1.0$ & various variances, as stated in Figure˜6: the histogram density plots compare the distribution of parameters for the attention layer L[12] in two contexts: the masking mechanism ($\bm{m}^{QKV}$) and the fine-tuned weighted product ($\bm{\theta}^{QKV}\odot\bm{m}^{QKV}$) for two tasks (Task 2 and Task 3). In Figure˜6(a), the $\bm{m}^{QKV}$ parameters are concentrated in a narrow range centered around 1. There is a slight variation between Task 2 and Task 3. However, the overall range remains consistent. $\bm{m}^{QKV}$ acts as a soft gating mechanism, controlling which pre-trained parameters are retained or modified during fine-tuning. In LABEL:{fig:soft_weight_cifar10}(b), the distribution of pre-trained weight ($\bm{\theta}^{QKV}$) is centered around 0, with a Gaussian-like spread. The distribution of soft-masked fine-tuned weight ($\bm{\theta}^{QKV}\odot\bm{m}^{QKV})$ remain similar to that of $\bm{\theta}^{QKV}$ but slightly narrower. The overlap between the two distributions indicates that the fine-tuning process maintains the general structure of the pre-trained weights while introducing task-specific modifications. By using soft-masks $\bm{m}^{QKV}$, Soft-TF ensures task-specific fine-tuning, reducing the risk of catastrophic forgetting and improving performance in various sequential tasks. The task-specific Soft-TF fine-tuning enables the models to infer task ID ideally in following Vision CIL scenarios.
Refer to caption Refer to caption
(a) $\bm{m}^{QKV}$ (b) $\bm{\theta}^{QKV}$ v.s. $\bm{\theta}^{QKV}\odot\bm{m}^{QKV}$
@@ 5.5 Soft-MSA Layers
Attention Layer. We compare attention maps from the 12th attention layer (L[12]) of two different models, DualPrompt and Soft-TF, applied to category 7 of the 10-Split-CIFAR100 dataset, as shown in Figure˜7. While DualPrompt distributes its attention more uniformly, Soft-TF prioritizes specific patches of the input image, likely those more relevant to Category 7. These attention maps highlight the differences in how DualPrompt and Soft-TF allocate attention in a Vision Transformer. Soft-TF’s sharper focus on essential features aligns with the reported superior performance on the 10-Split-CIFAR100 dataset. This visualization emphasizes the interpretability and task-specific optimization of Soft-TF over the more generalized behavior of DualPrompt.
@@ 6 Conclusion
Inspired by the Well-initialized Lottery Ticket Hypothesis (WLTH), we introduce Soft-Transformers (Soft-TF), a parameter-efficient framework for continual learning that leverages soft, real-valued subnetworks over a frozen pre-trained Transformer. Instead of relying on manually designed prompts or adapters, Soft-TF learns task-specific multiplicative masks applied to the key, query, value, and output projections in self-attention. These masks enable smooth and stable task adaptation while preserving shared representations. Combined with a lightweight dual-prompt mechanism, Soft-TF maintains strong knowledge retention and mitigates Catastrophic Forgetting (CF). Across multiple continual learning benchmarks, Soft-TF achieves state-of-the-art performance, consistently outperforming prompt-based, adapter-based, and LoRA-style baselines while requiring minimal additional parameters.
@@ A.1 Analysis of Soft-TransFormers (Soft-TF)
where $\bm{w}_{m}\neq\bm{w}_{q}$ since $\bm{m}$ is learnable parameters in Soft-TransFormers. Specifically, we could assume that $\bm{w}_{m}=(\bm{w}^{Q}\odot\bm{m}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{m}^{K})^{T}$ and $\bm{w}_{p}=(\bm{w}^{Q}\odot\bm{1}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{1}^{K})^{T}$ of LABEL:eq:lr_params.
Theorem A.2.
@@
Refer to caption Refer to caption
(a) L2P-PGP v.s. L2P+Soft-TF (b) DualPrompt v.s. DualPrompt+Soft-TF
Figure 8: Layer-wise(L[$\ast$]) Performances of Soft-TF on 10-Split-CIFAR100. Note that L[10,11,12] denotes Soft-TransFormer of 10, 11, 12 Layers.
Proof.
Using algebraic manipulations (completing the square), we obtain:
@@
ViT-B/16 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
BiC 5,000 - 81.42 17.31 73.02 6.23 64.63 22.25
DER++ 5,000 - 83.94 14.55 - - 66.73 20.67
iCaRL 5,000 - 66.00 5.33 78.02 5.80 - -
DER+MCG 2,000 - 67.62 14.64 65.84 13.72 - -
BiC 1,000 - 66.11 35.24 63.12 21.89 52.14 36.70
DER++ 1,000 - 61.06 39.87 - - 55.47 34.64
iCaRL 1,000 - 61.25 14.19 71.32 15.98 - -
FT - - 33.61 86.87 33.52 53.69 28.87 63.80
EWC - - 47.01 33.27 36.73 35.19 35.00 56.16
LWF - - 60.69 27.77 39.12 57.91 38.54 52.37
+ Soft-TF-L[3,4,5] 6.93M Prompt ID 91.77 / 3.37 13.87K / 80 94.43 / 2.02 15.60K / 104 74.70 / 6.46 15.35K / 52
+ Soft-TF-L[3,4,5] 6.93M Gradient ID 93.76 / 1.83 13.87K / 130 95.38 / 1.73 15.60K / 163 82.15 / 2.20 15.35K / 82
+ Soft-TF-L[10,11,12] 6.93M Gradient ID 97.87 / 0.21 13.87K / 130 99.05 / 0.24 15.60K / 163 82.38 / 0.59 15.35K / 82
Upper-Bound of DualPrompt 0.03M - 90.85 - 90.85 - 79.13 -
Upper-Bound of Soft-TF 6.93M - 95.90 - 95.90 - 80.21 -
In terms of Soft-TransFormers $\bm{w}_{m}=(\bm{w}^{Q}\odot\bm{m}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{m}^{K})^{T}$, we have
@@
A.2 Experimental Details
For fair comparisons with the baselines (Wang et al., 2022d, c; Qiao et al., 2024), we use ViT B/16 (Dosovitskiy et al., 2021) pre-trained on ImageNet-21K as our image encoder, which is kept frozen during training. We train and test on a single Quadro RTX 8000-48GB GPU for baselines and our Soft-TransFormers with Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$.
We adhere to the experimental settings outlined by (Qiao et al., 2024) to validate our method’s effectiveness. When comparing our approach with L2P-PGP and Soft-Transformer on the 10/20-Split-CIFAR100 and 10-Split-TinyImageNet datasets, we train the network for 5 epochs with a batch size of 16 and set the prompt length to 5. For the 10-Split-ImageNet-R dataset, we use 50 epochs, a batch size of 16, and a prompt length of 30. In comparison with DualPrompt-PGP and Soft-TransFormers on the 10/20-Split-CIFAR100 dataset, we train the network for 20 epochs with a batch size of 24 and set the expert prompt length to 5. For the 10-Split-TinyImageNet dataset, we use 5 epochs, a batch size of 24, and an expert prompt length of 5. For the 10-Split-ImageNet-R dataset, we set the epochs to 50, the batch size to 24, and the expert prompt length to 20. Additionally, in all benchmark data sets, the general prompt length is set to 5, and the location inserted into the prompt is kept consistent.
Table 6: Performances of Class Incremental Learning (CIL) in terms of accuracy and forgetting on $10/20$-Split-CIFAR100 and 10-Split-ImageNet-R.
ViT-B/16 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
S-Prompt 0.03M Prompt ID 87.57 / 3.63 12.12K / 76 84.90 / 7.05 11.60K / 78 74.25 / 4.73 13.10K / 47
CODA-Prompt 0.03M Prompt ID 86.94 / 4.04 12.12K / 76 84.70 / 7.11 11.60K / 78 74.26 / 5.17 13.10K / 47
HiDe-Prompt 0.03M Prompt ID 92.61 / 1.50 12.12K / 76 94.28 / 1.78 11.60K / 78 75.06 / 4.09 13.10K / 47
NoRGa 0.03M Prompt ID 94.48 / 1.44 12.12K / 76 97.39 / 0.58 11.60K / 78 75.40 / 4.59 13.10K / 47
+ Soft-TF-L[10,11,12] 6.93M Prompt ID 98.25 / 0.31 13.87K / 80 98.30 / 0.30 15.60K / 104 91.94 / 0.34 15.35K / 52
Comparisons with Prompt-based Baselines. We evaluate the performance of Soft-TF against representative prompt-based continual learning methods, including L2P (Wang et al., 2022d), DualPrompt (Wang et al., 2022c), CODA-Prompt (Smith et al., 2023b), S-Prompt (Wang et al., 2022a), HiDe-Prompt (Wang et al., 2023), and NoRGa (Le et al., 2024), as summarized in Table˜6. While these approaches demonstrate that improved prompt representations generally lead to better continual learning performance, their adaptation capacity remains fundamentally constrained by prompt-level conditioning. In contrast, Soft-TF consistently achieves state-of-the-art results by directly modulating attention projections through task-specific soft subnetworks, providing substantially greater representational expressiveness than prompts alone. These results indicate that, even when compared to strong and diverse prompt-based baselines, Soft-TF offers a more effective and scalable mechanism for continual adaptation.